i need to solve for x in these equations.Verify th esolution.
1. 28+x=7-2x
2. 2x=25+3x
3. 5(x-3)+2(3-x)=12
4. 18-(x+3)-2x=4+(x-5)

Answer by unlockmath(1688)

Hello,
I'll help you on the first one--let's see if you can figure out the others.
28+x=7-2x
First subtract x from both sides to get:
28=7-3x
Now subtract 7 from both sides to get:
21=-3x
Now divide -3 into both sides to result in:
x=-7
We can check to see if this is correct by plugging the answer in.
(Hint: What we try to do in these equations is get the variable, in this case "x" all by itself.
Make sense?
RJ
www.math-unlock.com